OpenStudy (anonymous):

f(x)= the square root of 6-x and g(x) = x+1 find (f+g)(15)

OpenStudy (anonymous):

f+g or fog??

OpenStudy (anonymous):

(f+g) ... the answer choices are a) 19 b)20 c)21 d)undefined

OpenStudy (anonymous):

\[(f+g)(15)=f(15)+g(15)\] does not exist because \[f(15)\] is not a real number

OpenStudy (anonymous):

if it is undefined i would still like to know how to go about it though... lost on where to start

OpenStudy (anonymous):

@Arnab09 do you agree ?

OpenStudy (anonymous):

replace x by 15

OpenStudy (anonymous):

\[f(15)=\sqrt{6-15}=\sqrt{-9}\]which is not a real number

OpenStudy (anonymous):

Thank you so much for explaining it !

OpenStudy (anonymous):

yw